# Pattern formation in a coupled driven diffusive system

### v1

@@ Appendix A Gradient expansion
Note that, for convenience, we collected all the linear/quadratic terms on the first/second line. From the GE, we find the functional forms contained in expression (III) for the linear coefficients listed here  [^19] For simplicity, we suppressed all the “physical units” for the coefficients, e.g., $D\propto\ell^{2}\tau^{-1}$, etc. Their detailed forms play no role in the numerical studies.. Determining the dependence of the nonlinear couplings on $\rho$ through the GE is nontrivial, as the expression (22) includes contributions of all orders in the fields $\rho_{A}(\mathbf{r},t)$ and $\rho_{B}(\mathbf{r},t)$, not just quadratic. For this reason, we choose to leave the quadratic coefficients as free parameters in the present study. Work is in progress to determine the exact functional dependence of the different nonlinear couplings on $\rho$ [48].
Appendix B Pseudospectral exponential time-differencing method
@@
The noise $\eta\equiv\eta(\mathbf{q},t)$ is a complex Gaussian variable, with zero mean and correlation
@@
The pseudospectral method is designed to numerically integrate partial differential equations, where we work directly with Eq. (25) in Fourier space to solve Eq. (III) by choosing an appropriate time-discretizing scheme [70]. The word pseudo refers to the fact that nonlinearities are calculated in direct space by means of the inverse Fourier transform, therefore we do not work fully in reciprocal, or Fourier, space. This method to calculate the nonlinearities introduces aliasing errors [71], which must be dealt with using dealiasing techniques, see Appendix C.
@@ Appendix C Dealiasing
$\rho_{\mathrm{sharp}}(k_{x},k_{y})=\begin{dcases}1\qquad\mathrm{if}\quad\frac{2k}{N}\leq\frac{2}{3}\\ 0\qquad\mathrm{if}\quad\frac{2k}{N}>\frac{2}{3}\end{dcases}$
$\rho_{\mathrm{smooth}}(k_{x},k_{y})=\exp\left[-36\left(\frac{2k}{N}\right)^{36}\right],$
where $N$ is the number of points in the lattice and $k=|\mathbf{k}|$ is the modulus of the mode index in Fourier space (see Appendix B). The effect of the first kernel is simply to abruptly set to zero $1/3$ of the amplitudes of the higher modes, while the second kernel sets the amplitudes of the higher modes to zero smoothly. Using the smooth dealiasing kernel, we keep about $12\sim 15\%$ more modes than the sharp dealiasing kernel. For $\rho_{\mathrm{smooth}}$, we use in the exponential the same factor of $36$ that is employed in [42].

### v2

@@ Appendix B Gradient expansion II – coefficients
The remaining seven nonlinear coefficients $\lambda_{\pm},\tilde{\lambda}_{\pm}$ and $g_{\pm,0}$ accompanying the quadratic terms in Eq.(25) are still undetermined. Determining their dependence on $\rho$ through the GE is nontrivial, as the expression (18) includes contributions of all orders in the fields $\rho_{A,B}$, not just quadratic. For this reason, we choose to leave the quadratic coefficients as free parameters in the present study. Work is in progress to determine the exact functional dependence of the different nonlinear couplings on $\rho$ [43].
This is a large parameter space (four $\lambda$’s, three $g$’s and $\sigma$), but a significant portion of it simply parametrizes the equilibrium $(\delta=0)$ behavior of the model. Therefore, we can introduce some simplifications to the equilibrium terms in order to better elucidate just the non-equilibrium features of interest. The specific simplifying approximations are
  1. 1.
  2. 2.
    Since stability for $\phi_{+}$ is always provided by $D_{+}>0$, we will drop the term $\lambda_{+}\phi^{2}_{+}$ from our equations  [^19] Note that such a term originates from a $\phi_{+}^{3}$ potential in the free-energy and, as such, is unstable. Therefore, one must include an additional $\gamma_{+}\,\phi^{3}_{+}$ (a $\phi_{+}^{4}$ potential) in Eq. (25) to ensure the stability of the theory. Since we avoid over-complicating our equations by seeking a minimal description and also considering that the DF is stable to linear-order in $\phi_{+}$, we can safely neglect the term proportional to $\lambda_{+}$.. This choice is additionally supported by a weakly nonlinear stability analysis  [^20] Based on the approach of Cross et al. [3, 5], the weakly nonlinear stability analysis here consists of assuming that a dominant role in the instability of the homogeneous regime is played by a particular mode of $\phi_{-}$ with the Ansatz $\phi_{-}(\mathbf{r},t)=\alpha(t)\,\mathrm{Re}\,\{e^{i\Psi(t)+i\mathbf{q}\cdot\mathbf{r}}\}$. As AA-$B$ exclusion tends to favor vacancies between the charged regions, it is reasonable to assume $\beta(t)\,\mathrm{Re}\,\{e^{i2\Psi(t)+i\mathbf{q}\cdot\mathbf{r}}\}$ for $\phi_{+}(\mathbf{r},t)$. In fact, the location of the peaks of the structure factors in the striped regime ($q_{-}=q^{\ast}=q_{+}/2$) displays this type of entraining of $\phi_{+}$ to $\phi_{-}$. Also, a perturbative scheme indicates $\beta\simeq\alpha^{2}$. Details are found in [43]. near the critical $\rho^{\mathrm{FT}}_{c}$. We find that $\phi_{+}$ is of the same order as $\phi_{-}^{2}$, so we drop terms with powers of $\phi_{+}$ larger than 1 when we study the high-density regime.
  3. 3.
It is also important to observe that, just as in mean-field theories of other statistical systems, different coarse-graining prescriptions (e.g., the present GE of the FLM or the MSRJD/Doi-Peliti developed in [2]) lead to different values for these coefficients, but they generally agree on their order of magnitude. Using the set of simplifying assumptions in Eq. (25) together with the linear coefficients in Eq. (B), we can write down the final form of the field equations used in this work: Eq. (III).
Appendix C Pseudospectral exponential time-differencing method
@@
The noise $\eta\equiv\eta(\mathbf{q},t)$ is a vector with complex Gaussian random variable components, with zero mean and correlation
@@
Note that, since $\Phi(\mathbf{r})$ is a real field, then $\tilde{\Phi}(\mathbf{q})$, the Fourier transform, satisfies the condition $\tilde{\Phi}(-\mathbf{q})=\tilde{\Phi}^{*}$. This should be true for the noise as well.
The pseudospectral method is designed to numerically integrate partial differential equations, where we work directly with Eq. (30) in Fourier space to solve Eq. (III) by choosing an appropriate time-discretizing scheme [67]. The word pseudo refers to the fact that nonlinearities are calculated in direct space by means of the inverse Fourier transform, therefore we do not work fully in reciprocal, or Fourier, space. This method to calculate the nonlinearities introduces aliasing errors [68], which must be dealt with using dealiasing techniques, see Appendix D.
@@ Appendix D Dealiasing
$\rho_{\mathrm{sharp}}(\mathbf{k})=\begin{dcases}1\qquad\mathrm{if}\quad\frac{2k}{N}\leq\frac{2}{3}\\ 0\qquad\mathrm{if}\quad\frac{2k}{N}>\frac{2}{3}\end{dcases}$
$\rho_{\mathrm{smooth}}(\mathbf{k})=\exp\left[-36\left(\frac{2k}{N}\right)^{36}\right],$
where $N$ is the number of points in the lattice and $k=|\mathbf{k}|$ is the modulus of the mode index in Fourier space (see Appendix C). The effect of the first kernel is simply to abruptly set to zero $1/3$ of the amplitudes of the higher modes, while the second kernel sets the amplitudes of the higher modes to zero smoothly. Using the smooth dealiasing kernel, we keep about $12\sim 15\%$ more modes than the sharp dealiasing kernel. For $\rho_{\mathrm{smooth}}$, we use in the exponential the same factor of $36$ that is employed in [37].